OpenStudy (anonymous):

You have a jar of pennies and quarters. You want to choose 15 coins that are worth exactly $4.35 A. Write and solve a system of equations that models the situation B. is your solution reasonable in terms of the original problem? explain.

OpenStudy (anonymous):

someone please help me

OpenStudy (snowsurf):

Lets have x be pennies and y be quarters Total coins will be x + y = 15 We know that penny is 1 cent so that will be 0.01 dollars and a quarter is 0.25 dollars. 0.01x + 0.25y = 4.35 So you have two equations x + y = 15 0.01x + 0.25y = 4.35 Just substitute one into the other. 0.01x + 0.25 (15-x) = 4.35 Solve for x and that will give you number of pennies then you can plug x into x + y =15 to get number of quarters.